The age of father is twice the age of elder son.Ten years hence,the age of father will be three times the younger son.If the difference in the ages of two sons is 15 years,Then the age of father is?
Answer by mananth(16946)

let elder son be x
father = 2x
y = younger son=(x-15)
10 years later
2x+10=3(x-15+10)
2x+10=3(x-5)
3x-2x=25
x=25
y=10
father = 50